If a cookie recipe to bake 36 cookies calls for 1/4 teaspoon of salt, how much salt would be required to bake 20 dozen cookies?
Answer by checkley71(8403)

36:1/4 AS 240:X
36X=240*1/4
36X=60
X=60/36
X=5/3
X=1 2/3 TEASPOONS OF SALT IS REQUIRED.